OpenStudy (anonymous):

Find f(x) if f'(x) = 7 x4 and f(1) = 4.

OpenStudy (anonymous):

can you rewrite that please? :D

OpenStudy (lgbasallote):

uhhh could you put the "f'(x) =" in one line? izkinda confusing..

OpenStudy (anonymous):

how do u rewrite it? lol do i close it then open it again?

OpenStudy (anonymous):

no just re-put your question on this thread

OpenStudy (lgbasallote):

or there is an Edit Question link beside Report Abuse

OpenStudy (lgbasallote):

on the question itself i mean..

OpenStudy (anonymous):

find f(x) if f '(x)= 7/x^4 and f(1)=4

OpenStudy (anonymous):

f'(x) = 7/x^4 \[\int\limits_{}^{} f'(x) = \int\limits_{}^{}7/x^4\] by the fundamental theorem of calculus \[\int\limits_{}^{} f'(x) = f(x)\] so: \[f(x) = \int\limits_{}^{}7/x^4\] 1/x^4 can be rewritten as x^-4, so: \[f(x) = \int\limits_{}^{} 7x ^{-4}\] integrating that, we get: f(x) = (-7/3)(x^-3) + c f(x) = (-7/3)(1/x^3) + C f(1) = 4 so plugging that in: 4 = (-7/3) (1/(1)^3) + C 4 = (-7/3) + C C = 19/3 our final answer should be: f(x) = (-7/3)(1/x^3) + 19/3 somebody check my work please!

OpenStudy (anonymous):

well the closest answer to yours in my answers is (-7/3) (x^-3) + 19/3 so idk if thats wat u meant

OpenStudy (anonymous):

but i appreciate your help!

OpenStudy (anonymous):

yes that is what i meant :D